Three years ago, Milcah is 3 times older than Joshua. 4 years from now, Milcah is 8 yrs. older than Joshua. Find Joshua's and Milcah's age.
Please answer, I need this urgent, I can't able to answer because I don't know how to solve age problem.. I really much appreciated your concerns.. Thank You!

Answer by TP(29)

Let m be Milcah's age now and j be Joshua's age now.
From what we are told about the ages 3 years ago we get the equation
m-3=3*(j-3) (i)
and from what we are told about the ages 4 years from now we get the equation
m+4=j+4+8, that is m+4=j+12. If we subtract 4 from each side of this equation we get that m=j+8. (ii)
Now we can replace m in equation (i) by j+8 so that equation (i) becomes
j+8-3=3*(j-3)
So j+5=3*(j-3)
Now multiply the bracket out and we get
j+5=3j-9
Now subtract j from both sides and we get
5=2j-9
Now add 9 to both sides and we get
14=2j
This means that j=7.
To find m replace j by 7 in equation(ii) and we get m=7+8 =15
So Milcah is 15 years old and Joshua is 7 years old.ANS
(NB I will leave it up to you to check the answers)